# Automated System-level Testing of Unmanned Aerial Systems

### v1

@@ Abstract
Unmanned aerial systems (UAS) rely on various avionics systems that are safety-critical and mission-critical. A major requirement of international safety standards is to perform rigorous system-level testing of avionics software systems. The current industrial practice is to manually create test scenarios, manually/automatically execute these scenarios using simulators, and manually evaluate outcomes. The test scenarios typically consist of setting certain flight or environment conditions and testing the system under test in these settings. The state-of-the-art approaches for this purpose also require manual test scenario development and evaluation. In this paper, we propose a novel approach to automate the system-level testing of the UAS. The proposed approach (AITester) utilizes model-based testing and artificial intelligence (AI) techniques to automatically generate, execute, and evaluate various test scenarios. The test scenarios are generated on the fly, i.e., during test execution based on the environmental context at runtime. The approach is supported by a toolset. We empirically evaluate the proposed approach on two core components of UAS, an autopilot system of an unmanned aerial vehicle (UAV) and cockpit display systems (CDS) of the ground control station (GCS). The results show that the AITester effectively generates test scenarios causing deviations from the expected behavior of the UAV autopilot and reveals potential flaws in the GCS-CDS.
@@
Model-based Testing, Artificial Intelligence, Deep Reinforcement Learning, Unmanned Aerial Systems, Testing Automation
@@ 1 Introduction
A traditional method of testing UAS at the system level includes a human tester (pilot) who knows UAV flight behavior and has expertise with UAV flight simulators. The tester is provided with the system under test (SUT) and the specifications of the SUT. Using this information, the tester manually designs test scenarios. For each test scenario, the tester simulates the behavior of a UAV in a particular environment with the help of a flight simulator, and at each flight operation, observes the behavior of the SUT by comparing the expected behavior mentioned in the specifications. Testing in this way usually requires several days and only a limited set of potential scenarios can be tested. The existing approaches available in the literature also require manual test scenario development and manual test evaluation (e.g., [3, 4, 5, 6]). The activity of manual testing is tedious and error-prone due to a large number of potential scenarios that need to be tested in the extremely dynamic environment of the UAV.
@@
In this paper, we propose a novel approach to automate the system-level testing of UAS. To automate the manual testing of UAS, a human-level control on a flight simulator and a mechanism to model the SUT specifications are required. Therefore, our approach (AITester) utilizes artificial intelligence (AI) and model-based testing (MBT) techniques to automatically generate, execute, and evaluate various test scenarios. We use deep reinforcement learning (an AI technique) to get human-level control on a flight simulator and MBT to model the specifications of SUT. The test scenarios are generated on the fly, i.e., during test execution based on the environmental context at runtime. For this purpose, we propose a UML profile for modeling the structural aspects of UAS SUT and a UML profile for modeling the flight behavior of the UAV. The structural profile is used to model the domain concepts of the UAS SUT and the flight behavior profile is used to model the abstract flight behavior of the UAV. Our approach requires the constraints (written in Object Constraint Language (OCL) [7]) specified on the UAS SUT domain model and state invariants on the flight behavior model. The UAS SUT domain model, UAV flight behavior model in the form of UML state machine, and the expected behavior of the SUT in the form of OCL constraints are used as input to our approach. Using the UAV flight behavior information, a deep reinforcement learning algorithm is trained to explore different flight test scenarios while interacting with the environment at runtime. The flight information received from the environment is used to determine the current state of the UAV. This information is also used to populate an instance model of the UAS SUT domain model. The instance model representing the current flight information of the UAS SUT is used to evaluate OCL constraints. The AITester gets rewards based on the correct selection of action and the number of OCL constraints violated based on that action. During the testing process, the AITester continues to explore different flight states, learns to violate OCL constraints (expected behavior of the SUT) based on actions, and compiles the results for further analysis by avionics testers.
@@
We develop a toolset that implements our proposed approach to support testing automation. We perform two experiments to evaluate AITester using two different subsystems of the UAS. For the first experiment, we used an autopilot of the UAV as a SUT. We use ArduCopter in the first experiment which is a widely-used open-source autopilot system. For the second experiment, we use the cockpit display systems (CDS) of the GCS as a SUT. We use an industrial case study of GCS-CDS for the second experiment. For UAV autopilot, the results show that AITester effectively generates test scenarios provoking deviations from the expected behavior of the UAV autopilot. For GCS-CDS, the results show that AITester is effective in finding four different types of faults and in exploring diverse test scenarios. Overall, the results of both experiments indicate that AITester effectively generates test scenarios causing deviations from the expected behavior of the SUT that lead to potential faults.
@@
  • We develop a toolset that implements our proposed approach to support testing automation and to enable further research and development. The toolset is publicly available at an online repository [^1] https://github.com/hassansartaj/uast-toolset.
  • We perform two experiments to evaluate AITester using two different subsystems of the UAS. The first experiment was performed using a UAV autopilot as a SUT, extending the pilot experiment with a new research question. The second experiment was performed using the CDS of GCS as a SUT to evaluate the applicability of the proposed approach to another UAS subsystem.
The remaining portion of the paper is organized as follows. Section 2 presents the background of unmanned aerial vehicles, model-based testing, reinforcement learning, and deep learning. Section 3 describes our proposed approach in detail. Section 4 provides a discussion on the toolset that implements the proposed approach. Section 5 provides the empirical evaluation of the proposed approach. Section 6 discusses the works related to this paper. Section 7 outlines limitations of proposed approach. Finally, Section 8 concludes the paper.
@@ 2.1 Unmanned Aerial Systems (UAS)
A GCS is a ground-based cockpit for the monitoring and control of a UAV. The GCS consists of the cockpit display systems (CDS), a flight control system, a mission planner, a communication channel, a navigation system, 2D/3D maps, and a payload monitoring and control system. The CDS of the GCS provides various types of interfaces with different flight instruments to show the UAV flight information [10]. These user interfaces display different types of information received from the different avionics systems of a UAV during the flight. For example, the navigation system generates information such as latitude, longitude, and altitude using sensors such as gyros, accelerometers, and GPS. This information is necessary for a UAV to accurately determine its current position, next waypoint to follow, and navigate on the planned route. The flight control system is used to generate control and navigational commands for the UAV. The mission planner supports UAV profile management, route planning, path optimization, payload setup, and planning unforeseen situations. The communication channel is responsible for supporting all types of communications between a UAV and a GCS. The navigation system is used to monitor the navigational information (such as latitude, longitude, and altitude) of the UAV. The 2D/3D maps support interactive mission planning, mission execution, and observing the state of the UAV during flight. The payload monitoring and control system helps an operator in observing the status of the payload and in controlling the payload.
@@ 2.2 Model-based Testing
Model-based testing (MBT) is a testing technique that uses models of the system under test (SUT) to support automation in various testing activities such as test case generation and test execution. For conventional system testing, where it is possible to construct complete models of expected behavior, MBT provides a comprehensive solution for automated test generation, test execution, and test evaluation [11]. The MBT process generally comprises five steps [12]. In the first step, the abstract models of the SUT are developed based on the requirements or specifications using any modeling language such as Unified Modeling Langauge (UML) [13]. The SUT models capture the system’s functionality in the form of various UML models (diagrams) such as a class diagram, a state machine, and a sequence diagram. Different UML models support the automation of several testing activities. For example, UML state machines support the generation of test sequences [10]. The constraints specified in the requirements are modeled using a constraint specification language, i.e., Object Constraint Language (OCL) [7, 14]. The SUT models developed in UML are augmented with the constraints written in OCL. The second step is to determine the criteria for test case selection. The criteria for selecting test cases can be defined based on requirements (e.g., requirement coverage) or using the models of SUT (e.g., state/transition/transition-pair coverage of the state machine). In the third step, the specifications of test cases are defined based on the criteria for test selection. In the fourth step, the models of SUT (created in the first step) and test case specifications (defined in the third step) are used for the generation of test cases. In the fifth step, the generated test cases are executed using the test execution platform that is suitable for the SUT. In the end, the test verdict is determined by comparing the test execution output with the SUT expected output. If the expected output matches the test execution output, the verdict is pass and fail otherwise.
@@ 2.3 Reinforcement Learning (RL)
Reinforcement learning (RL) is an artificial intelligence technique in which an agent learns based on its interaction with the environment. The RL agent, in a particular state, performs some action in the environment. Based on the action, the environment returns the next state and the reward. Therefore, the RL problem can be formulated as a Markov Decision Process (MDP). Fig. 1 shows an MDP in which an agent in a state $S_{t}$ with reward $R_{t}$ performs an action $A_{t}$ on the environment at time t. Based on the action, the environment returns the next state $S_{t+1}$ and the reward $R_{t+1}$. According to the reward received from the environment, the agent learns whether the action was right or wrong. The process of learning continues until the agent learns an optimal policy that can maximize the expected reward to explore the environment. RL is used for multiple applications in a variety of different domains [16] such as robotics, transportation, and games. In the context of UAVs, it has been used for various purposes such as for the UAV attitude control [17] and the robust control of the autonomous helicopter [18].
@@ 2.4 Deep Learning
It is a machine learning method that simulates a part of the human brain. For this purpose, it uses artificial neural networks (ANNs) inspired by the biological neural networks of the human brain [19]. The ANN can be a single-layer network with only input and output layers. It can be a multi-layer network with an input layer, an output layer, and several hidden layers in the middle. Multi-layer neural networks are also called deep neural networks. There are three commonly used types of ANNs, (i) feed-forward neural network, (ii) convolutional neural network (CNN), and (iii) recurrent neural network (RNN). In a feed-forward neural network, the information from the input layer flows toward the output layer after performing the function approximations in the intermediate levels. The feed-forward neural networks are also known as deep feed-forward networks or multi-layer perceptrons. In RNN, the information from the input layer flows toward the output layer and can be fed back to the neurons in the previous layer. In CNN, the information from the input layer flows toward the output layer by performing different convolution operations in the middle.
@@ 3 Approach
In this section, we discuss the proposed approach for the automated system-level testing of the UAS. The proposed approach is developed based on the idea presented in previous work [8]. First, we give a brief overview of the proposed approach. Second, we describe the modeling methodology including a discussion of the proposed profiles. In the end, we discuss the deep reinforcement learning-based approach for the testing of a UAS SUT based on environmental context at runtime.
@@ 3.2.3 Modeling Guidelines
The flight behavior model of the UAV needs to be modeled using the UML state machine. For this purpose, all flight states and events need to be modeled using the states and events stereotypes from the UAV behavioral profile. The avionics tester is required to model the end-to-end flight behavior of the UAV, i.e., to take off from the ground, perform different flight operations, and land on the ground. For example, it is not possible for a UAV to fly in the air without initial ground operations, therefore, the UAV flight state machine without ground states/events will not be useful. All the state machine events are required to be modeled as UML CallEvents. Other types of events such as SignalEvent, TimeEvent, and ChangeEvent are simulator-specific events that can not be controlled using the DRL algorithm. In the case of guard conditions, our approach requires modeling the guards in the form of OCL constraints or invariants. While modeling transitions, the avionics tester needs to model all possible transitions and events for a flight state.
@@ 3.3.1 UAS SUT Domain Modeling
The domain model of the UAS SUT serves two purposes. First, it is used to provide the UAS SUT domain classes and properties to the environment used for the reinforcement learning algorithm (Section 3.4). This information becomes the observation space of the environment that is observed during the UAV flight. The second use of the domain model is to evaluate OCL constraints. For this purpose, an instance of the UAS SUT domain model is created and the slots are filled using the UAV flight data. The OCL constraints modeled for the corresponding flight state are evaluated on the instance model. The evaluation results are used later on for the reward calculation (Section 3.4.2).
@@ 3.3.2 Flight Behavior Modeling
The flight behavior of the UAV is represented by many abstract flight states and events. Fig. 7 and Fig. 8 show the flight behavior modeled as a UML state machine for the ArduCopter. The flight state machine contains various abstract flight states modeled as instances of meta-states from the UAV behavioral profile. Similarly, all the events are modeled as instances of various flight events from the UAV behavioral profile.
@@
Ascend is the state in which the ArduCopter is making a climb i.e., increasing the altitude. Descend is the state in which the ArduCopter is making a descending i.e., decreasing the altitude. The Descend state also has a sub-state machine for turns because the ArduCopter can move straight or it can take a left turn or right turn. The Loiter state represents a flight phase in which the ArduCopter is flying in a small circle. During the Loiter state, the ArduCopter can loiter in a clockwise or counterclockwise direction. Thus, it has a sub-state machine for right and left turns for loitering in a clockwise or counterclockwise direction. The state PositionHold corresponds to the flight phase in which the ArduCopter is holding its position while keeping the attitude constant. Therefore, PositionHold state does not have a sub-state machine for different types of turns. Similarly, the AltitudeHold state represents a flight phase in which the ArduCopter is flying at the same level while keeping the altitude constant. For the ArduCopter types of UAV, it is possible to maintain constant altitude and can move forward or backward or turn left or right.
Listing 1: Some of the OCL constraints including the general and the state invariants
--altitude values range (in meters)
C1: context ArduCopter inv: self.location.altitude_AGL>0 and self.location.altitude_AGL<=300
--distance values range (in meters)
C2: context ArduCopter inv: self.rangefinder.distance>0 and self.rangefinder.distance<=5000
--State invariants for Takeoff
C3: context ArduCopter inv: self.oclIsInState(Takeoff) and self.thrust>0 and self.thrust<1
C4: context ArduCopter inv: self.oclIsInState(Takeoff) and self.location.altitude_AGL>0 and self.location.altitude_AGL<=50
--State invariants for Landing
C5: context ArduCopter inv: self.oclIsInState(Landing) and self.airspeed>0 and self.airspeed<=5
C6: context ArduCopter inv: self.oclIsInState(Landing) and self.speed.vz>0 and self.speed.vz<=5
--State invariants for Ascend
C7: context ArduCopter inv: self.oclIsInState(Ascend) and self.airspeed>0 and self.airspeed<100
C8: context ArduCopter inv: self.oclIsInState(Ascend) and self.groundspeed>0 and self.groundspeed<10
--State invariants for Descent
C9: context ArduCopter inv: self.oclIsInState(Descent) and self.airspeed>0 and self.airspeed<100
C10: context ArduCopter inv: self.oclIsInState(Descent) and self.location.altitude_AGL>10 and self.location.altitude_AGL<100

### v2

@@ Abstract
Unmanned aerial systems (UAS) rely on various avionics systems that are safety-critical and mission-critical. A major requirement of international safety standards is to perform rigorous system-level testing of avionics software systems. The current industrial practice is to manually create test scenarios, manually/automatically execute these scenarios using simulators, and manually evaluate outcomes. The test scenarios typically consist of setting certain flight or environment conditions and testing the system under test in these settings. The state-of-the-art approaches for this purpose also require manual test scenario development and evaluation. In this paper, we propose a novel approach to automate the system-level testing of the UAS. The proposed approach (namely AITester) utilizes model-based testing and artificial intelligence (AI) techniques to automatically generate, execute, and evaluate various test scenarios. The test scenarios are generated on the fly, i.e., during test execution based on the environmental context at runtime. The approach is supported by a toolset. We empirically evaluated the proposed approach on two core components of UAS, an autopilot system of an unmanned aerial vehicle (UAV) and cockpit display systems (CDS) of the ground control station (GCS). The results show that the AITester effectively generates test scenarios causing deviations from the expected behavior of the UAV autopilot and reveals potential flaws in the GCS-CDS.
@@
Artificial Intelligence, Deep Reinforcement Learning, Unmanned Aerial Systems, UAV, Drones, GCS, Testing Automation
@@ 1 Introduction
A traditional method of testing UAS at the system level includes a human tester (pilot) who knows UAV flight behavior and has expertise with UAV flight simulators. The tester is provided with the system under test (SUT) and the specifications of the SUT. Using this information, the tester manually designs test scenarios. For each test scenario, the tester simulates the behavior of a UAV in a particular environment with the help of a flight simulator, and at each flight operation, observes the behavior of the SUT by comparing the expected behavior mentioned in the specifications. Testing this way usually requires several days and only a limited set of potential scenarios can be tested. The existing approaches available in the literature also require manual test scenario development and manual test evaluation (e.g., [3, 4, 5, 6]). The activity of manual testing is tedious and error-prone due to numerous potential scenarios that need to be tested in the extremely dynamic environment of the UAV.
@@
In this paper, we propose a novel approach to automate the system-level testing of UAS. To automate the manual testing of UAS, a human-level control on a flight simulator and a mechanism to model the SUT specifications are required. Therefore, our approach (namely AITester) utilizes artificial intelligence (AI) and model-based testing (MBT) techniques to automatically generate, execute, and evaluate various test scenarios. We use deep reinforcement learning (an AI technique) to get human-level control on a flight simulator and MBT to model the specifications of SUT. The test scenarios are generated on the fly, i.e., during test execution based on the environmental context at runtime. For this purpose, we propose a UML profile for modeling the structural aspects of UAS SUT and a UML profile for modeling the flight behavior of the UAV. The structural profile is used to model the domain concepts of the UAS SUT and the flight behavior profile is used to model the abstract flight behavior of the UAV. Our approach requires the constraints (written in Object Constraint Language (OCL) [7]) specified on the UAS SUT domain model and state invariants on the flight behavior model. The UAS SUT domain model, the UAV flight behavior model in the form of a UML state machine, and the expected behavior of the SUT in the form of OCL constraints are used as input to our approach. Using the UAV flight behavior information, a deep reinforcement learning algorithm is trained to explore different flight test scenarios while interacting with the environment at runtime. The flight information received from the environment is used to determine the current state of the UAV. This information is also used to populate an instance model of the UAS SUT domain model. The instance model representing the current flight information of the UAS SUT is used to evaluate OCL constraints. The AITester gets rewards based on the correct selection of action and the number of OCL constraints violated based on that action. During the testing process, the AITester continues to explore different flight states, learns to violate OCL constraints (expected behavior of the SUT) based on actions, and compiles the results for further analysis by avionics testers.
@@
We developed a toolset that implements our proposed approach to support testing automation. We performed two experiments to evaluate AITester using two different subsystems of the UAS. For the first experiment, we used an autopilot of the UAV as a SUT, specifically employing the ArduCopter, a widely-used open-source autopilot system. For the second experiment, we used the cockpit display systems (CDS) of the GCS as the SUT, particularly utilizing an industrial case study of GCS-CDS. For UAV autopilot, the results indicated that AITester effectively generates test scenarios provoking deviations from the expected behavior of the UAV autopilot. For GCS-CDS, the results showed that AITester is effective in finding four different types of faults and in exploring diverse test scenarios. Overall, the results of both experiments demonstrated that AITester effectively generates test scenarios causing deviations from the expected behavior of the SUT that lead to potential faults.
@@
  • We developed a toolset that implements our proposed approach to support testing automation and to enable further research and development. The toolset is publicly available at an online repository [^1] https://github.com/hassansartaj/uast-toolset.
  • We performed two experiments to evaluate AITester using two different subsystems of the UAS. The first experiment was performed using a UAV autopilot as a SUT, extending the pilot experiment with a new research question. The second experiment was performed using the CDS of GCS as a SUT to evaluate the applicability of the proposed approach to another UAS subsystem.
The remaining portion of the paper is organized as follows. Section 2 presents the background of unmanned aerial vehicles, model-based testing, reinforcement learning, and deep learning. Section 3 describes our proposed approach in detail. Section 4 provides a discussion on the toolset that implements the proposed approach. Section 5 provides the empirical evaluation of the proposed approach. Section 6 discusses the works related to this paper. Section 7 outlines the limitations of the proposed approach. Finally, Section 8 concludes the paper.
@@ 2.1 Unmanned Aerial Systems (UAS)
A GCS is a ground-based cockpit for monitoring and controlling a UAV. The GCS consists of the cockpit display systems (CDS), a flight control system, a mission planner, a communication channel, a navigation system, 2D/3D maps, and a payload monitoring and control system. The CDS of the GCS provides various types of interfaces with different flight instruments to show the UAV flight information [10]. These user interfaces display different types of information received from the different avionics systems of a UAV during the flight. For example, the navigation system generates information such as latitude, longitude, and altitude using sensors such as gyros, accelerometers, and GPS. This information is necessary for a UAV to accurately determine its current position, next waypoint to follow, and navigate on the planned route. The flight control system is used to generate control and navigational commands for the UAV. The mission planner supports UAV profile management, route planning, path optimization, payload setup, and planning unforeseen situations. The communication channel is responsible for supporting all types of communications between a UAV and a GCS. The navigation system is used to monitor the navigational information (such as latitude, longitude, and altitude) of the UAV. The 2D/3D maps support interactive mission planning, mission execution, and observing the state of the UAV during flight. The payload monitoring and control system helps an operator in observing the status of the payload and in controlling the payload.
@@ 2.2 Model-based Testing
Model-based testing (MBT) is a testing technique that uses models of the system under test (SUT) to support automation in various testing activities such as test case generation and test execution. For conventional system testing, where it is possible to construct complete models of expected behavior, MBT provides a comprehensive solution for automated test generation, test execution, and test evaluation [11]. The MBT process generally comprises five steps [12]. In the first step, the abstract models of the SUT are developed based on the requirements or specifications using any modeling language such as Unified Modeling Langauge (UML) [13]. The SUT models capture the system’s functionality in the form of various UML models (diagrams) such as a class diagram, a state machine, and a sequence diagram. Different UML models support the automation of several testing activities. For example, UML state machines support the generation of test sequences [10]. The constraints specified in the requirements are modeled using a constraint specification language, i.e., Object Constraint Language (OCL) [7, 14]. The SUT models developed in UML are augmented with the constraints written in OCL. The second step is to determine the criteria for test case selection. The criteria for selecting test cases can be defined based on requirements (e.g., requirement coverage) or using the models of SUT (e.g., state/transition/transition-pair coverage of the state machine). In the third step, the specifications of test cases are defined based on the criteria for test selection. In the fourth step, the models of SUT (created in the first step) and test case specifications (defined in the third step) are used for generating test cases. In the fifth step, the generated test cases are executed using the test execution platform that is suitable for the SUT. In the end, the test verdict is determined by comparing the test execution output with the SUT expected output. If the expected output matches the test execution output, the verdict is pass and fail otherwise.
@@ 2.3 Reinforcement Learning (RL)
Reinforcement learning (RL) is an artificial intelligence technique in which an agent learns based on its interaction with the environment. The RL agent, in a particular state, performs some action in the environment. Based on the action, the environment returns the next state and the reward. Therefore, the RL problem can be formulated as a Markov Decision Process (MDP). Fig. 1 shows an MDP in which an agent in a state $S_{t}$ with reward $R_{t}$ performs an action $A_{t}$ on the environment at time t. Based on the action, the environment returns the next state $S_{t+1}$ and the reward $R_{t+1}$. According to the reward received from the environment, the agent learns whether the action was right or wrong. The learning process continues until the agent learns an optimal policy that can maximize the expected reward to explore the environment. RL is used for multiple applications in a variety of different domains [16] such as robotics, transportation, and games. In the context of UAVs, it has been used for various purposes such as for the UAV attitude control [17] and the robust control of the autonomous helicopter [18].
@@ 2.4 Deep Learning
It is an AI technique that simulates a part of the human brain. For this purpose, it uses artificial neural networks (ANNs) inspired by the biological neural networks of the human brain [19]. The ANN can be a single-layer network with only input and output layers. It can be a multi-layer network with an input layer, an output layer, and several hidden layers in the middle. Multi-layer neural networks are also called deep neural networks. There are three commonly used types of ANNs, (i) feed-forward neural network, (ii) convolutional neural network (CNN), and (iii) recurrent neural network (RNN). In a feed-forward neural network, the information from the input layer flows toward the output layer after performing the function approximations in the intermediate levels. The feed-forward neural networks are also known as deep feed-forward networks or multi-layer perceptrons. In RNN, the information from the input layer flows toward the output layer and can be fed back to the neurons in the previous layer. In CNN, the information from the input layer flows toward the output layer by performing different convolution operations in the middle.
Long Short-Term Memory (LSTM). An LSTM network is a variant of RNNs, specifically designed to handle temporal data and capture the state context. It comprises one or more units, each containing input gates, output gates, forget gates, and memory cells. The core purpose of the memory cell is to capture the correlation between the current data at time t with preceding data points, i.e., (t-1)!. This mechanism allows LSTM networks to learn longer sequences, making them especially useful for tasks involving sequential data such as time-series analysis.
@@ 3 Approach
In this section, we discuss the proposed approach for the automated system-level testing of the UAS. The proposed approach is developed based on the idea presented in previous work [8]. First, we give a brief overview of the proposed approach. Second, we describe the modeling methodology including a discussion of the proposed profiles. In the end, we discuss the deep reinforcement learning technique employed in our approach for testing a UAS SUT based on environmental context at runtime.
@@ 3.2.3 Modeling Guidelines
The flight behavior model of the UAV needs to be modeled using the UML state machine. For this purpose, all flight states and events need to be modeled using the states and events stereotypes from the UAV behavioral profile. The avionics tester is required to model the end-to-end flight behavior of the UAV, i.e., to take off from the ground, perform different flight operations, and land on the ground. For example, a UAV cannot fly in the air without initial ground operations, therefore, the UAV flight state machine without ground states/events will not be useful. All the state machine events are required to be modeled as UML CallEvents. Other types of events such as SignalEvent, TimeEvent, and ChangeEvent are simulator-specific events that can not be controlled using the DRL algorithm. In the case of guard conditions, our approach requires modeling the guards in the form of OCL constraints or invariants. While modeling transitions, the avionics tester needs to model all possible transitions and events for a flight state.
@@ 3.3.1 UAS SUT Domain Modeling
The domain model of the UAS SUT serves two purposes. First, it is used to provide the UAS SUT domain classes and properties to the environment utilized for the reinforcement learning algorithm (Section 3.4). This information becomes the observation space of the environment that is observed during the UAV flight. The second use of the domain model is to evaluate OCL constraints. For this purpose, an instance of the UAS SUT domain model is created and the slots are filled using the UAV flight data. The OCL constraints modeled for the corresponding flight state are evaluated on the instance model. The evaluation results are used later for the reward calculation (Section 3.4.2).
@@ 3.3.2 Flight Behavior Modeling
The flight behavior of the UAV needs to be represented using abstract flight states and events. Fig. 7 and Fig. 8 show the flight behavior modeled as a UML state machine for the ArduCopter. The flight state machine contains various abstract flight states modeled as instances of meta-states from the UAV behavioral profile. Similarly, all the events are modeled as instances of various flight events from the UAV behavioral profile.
@@
Ascend is the state in which the ArduCopter is making a climb i.e., increasing the altitude. Descend is the state in which the ArduCopter is making a descending i.e., decreasing the altitude. The Descend state also has a sub-state machine for turns because the ArduCopter can move straight or it can take a left turn or right turn. The Loiter state represents a flight phase in which the ArduCopter is flying in a small circle. During the Loiter state, the ArduCopter can loiter in a clockwise or counterclockwise direction. Thus, it has a sub-state machine for right and left turns for loitering in a clockwise or counterclockwise direction. The state PositionHold corresponds to the flight phase in which the ArduCopter is holding its position while keeping the attitude constant. Therefore, PositionHold state does not have a sub-state machine for different types of turns. Similarly, the AltitudeHold state represents a flight phase in which the ArduCopter is flying at the same level while keeping the altitude constant. For the ArduCopter types of UAV, it is possible to maintain constant altitude and move forward or backward or turn left or right.